A rectangular garden is 4 feet by 3 feet a landscaper plants flowers 1 ft apart along the edges and corners how many plants does the landscaper need

Answer by ikleyn(52835)

.

      5 + 5 = 10 flowers along the two longer edges

PLUS

      2 + 2 = 4  flowers along the two shorter edges  ( ! the corners are just occupied !).


In all,  10 + 4 = 14 flowers.


ANOTHER  SOLUTION  is  THIS :

The perimeter is  4 + 3 + 4 + 3 = 14 ft.


Planting the flowers along the perimeter at 1 ft apart, you will plant 14 flowers.